# Two-Step QAOA: Enhancing Quantum Optimization by Decomposing One-Hot Constraints in QUBO Formulations

### v1

@@ Abstract
The Quantum Approximate Optimization Algorithm (QAOA) has shown promise in solving combinatorial optimization problems by leveraging quantum computational power. We propose a simple approach, the Two-Step QAOA, which aims to improve the effectiveness of QAOA by decomposing problems with one-hot encoding QUBO (Quadratic Unconstrained Binary Optimization) formulations. By identifying and separating the problem into two stages, we transform soft constraints into hard constraints, simplifying the generation of initial conditions and enabling more efficient optimization. The method is particularly beneficial for tackling complex societal problems that often involve intricate constraint structures.
@@ 1 Introduction
Quantum combinatorial optimization problems using quantum computers have seen various methods proposed for both problem formulation and actual solution implementation through the use of QAOA (Quantum Approximate Optimization Algorithm) [1]. QAOA can convert the formulation into operations called quantum gates according to theory and implement them on the quantum circuits of actual quantum computers or corresponding simulators. Currently, many societal problems worldwide are being tackled using algorithms like QAOA and quantum annealing, with applications expected in diverse fields such as logistics, finance, and natural language processing [2]. However, the performance of QAOA is highly dependent on the problem formulation. When dealing with QUBO (Quadratic Unconstrained Binary Optimization) formulations for societal problems [3], the relationship between constraint conditions and cost functions often leads to unnecessary parameter adjustments and computational load. Efficient computation necessitates innovations in formulation and subsequent embedding into quantum circuits.
When introducing penalty terms in QUBO, the adjustment variables that determine the influence of constraint conditions are introduced, making adjustments difficult depending on the state of the formulation. To eliminate the uncertainties associated with such soft constraints, the use of the Quantum Alternating Operator Ansatz (QAOA) is effective. This approach treats some soft constraints as hard constraints and solves these challenges by leveraging quantum entanglement and the mixer Hamiltonian.
In this paper, we propose a simple approach called "Two-Step QAOA." This approach enhances the effectiveness of QAOA by decomposing the QUBO formulation for constraints known as one-hot constraints and utilizing the Quantum Alternating Operator Ansatz. QUBO problems are widely used to generalize various combinatorial optimization problems and have been extensively studied due to their flexibility and expressiveness. However, QUBO problems often mix soft and hard constraints, complicating the adjustment process and making optimization challenging.
Two-Step QAOA divides specific problems into two quantum circuit stages, transforming soft constraints into hard constraints, simplifying the creation of initial conditions, and enabling more efficient optimization. In the first stage, we distinguish between soft constraints and cost functions from the structure of the QUBO formulation, handling the soft constraints separately and preparing the process to transform them into hard constraints and corresponding quantum states. In the second stage, we use the hard constraints transformed from the soft constraints to create a new QUBO formulation, optimizing it more effectively using QAOA.
@@
The objective of this study is to use the Two-Step QAOA approach to reduce the complexity of setting initial conditions and enhance the overall performance of QAOA. This is expected to open new avenues for solving complex societal problems such as logistics, resource allocation, and scheduling. Additionally, since the quantum state generation process is heuristic, various improvements, including approximations, can be anticipated.
@@ 3.2 Initialization
In a basic explanation of QAOA, we first assume a simple state, preparing all qubits in a uniform superposition. This state means that all qubits are equally likely to be in the state 0 or 1. The initial state is represented as follows:
$|\psi_{0}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{z\in\{0,1\}^{n}}|z\rangle$ (2)
@@ 3.3 Application of Cost and Mixer Hamiltonians
3.4 Parameter Optimization
The application of the cost and mixer Hamiltonians requires time parameters $\gamma$ and $\beta$. These parameters are optimized using classical optimization methods to maximize the probability of reaching the optimal solution.
3.5 Measurement and Result Extraction
The final state is measured to obtain the solution to the optimization problem. The measurement result is a bit string that represents the problem’s optimal solution.
3.6 Mathematical Representation
The overall process of QAOA is mathematically represented as follows. Starting from the initial state, the cost and mixer Hamiltonians are applied $p$ times:
@@ 4.1 Mixer Hamiltonian
In standard QAOA, the mixer Hamiltonian is typically the $X$ operator. However, in the Quantum Alternating Operator Ansatz, the mixer Hamiltonian is constructed using Pauli operators. This allows for explorations that cannot be achieved with just the $X$ Hamiltonian.
@@ 4.2 Initialization
Instead of assuming a simple state, the initial state in the Quantum Alternating Operator Ansatz is prepared as an eigenstate with complex quantum entanglement corresponding to the mixer Hamiltonian.
5 Cost and Constraint Terms in QAOA formulation
@@
In QAOA, constructing the cost Hamiltonian involves creating terms that minimize the objective function and satisfy necessary constraints. These terms are balanced using tuning parameters, presenting several challenges during the process.
5.1 Formulating Cost and Constraint Terms
The cost Hamiltonian $H_{C}$ is composed of a cost term $C$, which represents the objective function, and a penalty term $P$, which represents the constraints that must be satisfied. These terms are connected using a tuning parameter $\lambda$:
$H_{C}=C+\lambda P$ (7)
5.2 Balancing the Terms
The primary challenge is to balance the cost and constraint terms to ensure that the solution minimizes the objective function while satisfying all constraints. If the tuning parameter $\lambda$ is too large, the solution might prioritize satisfying the constraints over minimizing the cost, leading to suboptimal solutions. Conversely, if $\lambda$ is too small, the constraints might be violated.
5.3 Iterative Tuning of Parameters
@@
Determining the appropriate value for $\lambda$ often requires iterative optimization. This process involves adjusting $\lambda$ iteratively until a satisfactory balance is achieved, making sure that the final solution both minimizes the cost and satisfies the constraints.
The search space expands with the addition of constraint terms, making it harder to find the global minimum.
6 Incorporating Hard Constraints into Mixer Hamiltonian in QAOA: A Two-Step Initialization Approach
In specific formulations, such as when the constraint conditions can be described using an XY operator in the mixer Hamiltonian, it is possible to relax the constraints from the cost function. These constraints are excluded from the cost Hamiltonian as hard constraints and instead incorporated into the mixer Hamiltonian. This approach is particularly effective for problems involving one-hot encoding, where the constraints can naturally be expressed using the XY mixer. By leveraging the symmetry and properties of the XY operator, one-hot encoded variables can be efficiently managed, ensuring that only one variable in a set is activated while the others remain off. Utilizing the mixer Hamiltonian necessitates the preparation of an initial state corresponding to it. Various methods have been proposed for this purpose, including the use of Dicke states. In this case, the initial state can be easily created using a two-step approach.
@@ 6.2 Construct Mixer Hamiltonian
In this problem, QAOA is executed twice, each time using a different mixer Hamiltonian.
@@
$H_{M_{2}}=H_{XY}=\sum_{i,j}XY_{ij}+\ldots$ (10)
where $XY_{ij}$ represents the XY interaction between qubits $i$ and $j$.
@@ 6.3 Modify Cost Hamiltonian
Exclude the identified constraints from the cost Hamiltonian $H_{C}$, treating them as hard constraints now embedded in the mixer Hamiltonian. The modified cost Hamiltonian will focus solely on the objective function.
$H_{C}=C$ (11)
6.4 Prepare Initial State
Prepare the initial state that corresponds to the mixer Hamiltonian. Various methods have been proposed for this purpose, including the use of Dicke states [5]. In this case, the initial state can be efficiently created using a two-step approach:
Start with an initial superposition state:
@@ 6.4 Prepare Initial State
6.5 First QAOA Execution
Run QAOA with the constraints as the cost function and $X$ as the mixer Hamiltonian to generate the intermediate quantum state:
6.6 Second QAOA Execution
Use the intermediate quantum state as the starting point and run QAOA with the objective function as the cost Hamiltonian and the XY mixer Hamiltonian:
Optimize the parameters $\gamma_{1},\gamma_{2},\beta_{1},\beta_{2}$ through classical optimization techniques to find the optimal solution.

### v2

@@ Abstract
The Quantum Approximate Optimization Algorithm (QAOA) has shown promise in solving combinatorial optimization problems by leveraging quantum computational power. We propose a simple approach, the Two-Step QAOA, which aims to improve the effectiveness of QAOA by decomposing problems with k-hot encoding QUBO (Quadratic Unconstrained Binary Optimization) formulations. By identifying and separating the problem into two stages, we transform soft constraints into hard constraints, simplifying the generation of initial conditions and enabling more efficient optimization. The method is particularly beneficial for tackling complex societal problems that often involve intricate constraint structures.
@@ 1 Introduction
In modern research and business environments, there is a growing focus on optimizing entire processes by improving the efficiency of intermediate steps, which can significantly reduce costs while achieving the same objectives. Among these, combinatorial optimization problems represent a critical field that has been extensively studied using classical computers. For instance, challenges such as optimizing logistics, resource allocation, and scheduling are essential for everyday decision-making and have received considerable attention.
Moreover, in the emerging field of quantum computing, which is expected to serve as a next-generation computational platform, research and development in this domain are highly active [1] [2]. Leveraging the unique properties of quantum computers offers the potential to tackle large-scale, complex problems that are difficult to solve using traditional methods. Against this backdrop, the Quantum Approximate Optimization Algorithm (QAOA) has gained recognition as a key algorithm for efficiently addressing combinatorial optimization problems.
@@
Quantum combinatorial optimization problems using quantum computers have seen various methods proposed for both problem formulation and actual solution implementation through the use of QAOA (Quantum Approximate Optimization Algorithm) [3]. QAOA can convert the formulation into operations called quantum gates according to theory and implement them on the quantum circuits of actual quantum computers or corresponding simulators. Currently, many societal problems [4] worldwide are being tackled using algorithms like QAOA and quantum annealing, with applications expected in diverse fields such as logistics, finance, and natural language processing [5]. However, the performance of QAOA is highly dependent on the problem formulation.
The formulation of QAOA utilizes a problem representation called QUBO (Quadratic Unconstrained Binary Optimization) [6], which transforms problems into a quadratic expression. QUBO leverages the binary values 0 and 1 of quantum bits to define the problem. This formulation primarily consists of constraint terms, which represent the conditions that must be satisfied, and cost functions, which aim to optimize efficiency. These components are designed as terms in the equation and are combined for use. In particular, constraint terms are often formulated as penalty terms, where violating the constraints results in higher costs.
A major challenge with the QUBO format lies in the relationship between constraint terms and the cost function, which often leads to unnecessary parameter tuning and increased computational load. Such tuning frequently varies depending on the specific problem, making it difficult to predict in advance and even more challenging due to its dependence on the state of the formulation. To achieve efficient computation, addressing these issues through innovations in formulation and subsequent embedding into quantum circuits is essential.
One potential solution to eliminate uncertainties associated with such constraints is the effective use of the Quantum Alternating Operator Ansatz (QAOA). The Quantum Alternating Operator Ansatz modifies the mixer Hamiltonian, which is the exploratory component used in standard QAOA, in conjunction with the initial quantum state. This allows certain soft constraints to be omitted, thereby eliminating the need for complex adjustments. While not all types of constraints can be addressed, it is possible to handle problems involving one-hot constraints, which are commonly encountered in optimization contexts, by treating them as hard constraints and removing them from the formulation.
In this study, we focus on the XY mixer, a mixer that swaps the values of quantum bits, as the mixer Hamiltonian. To use this mixer, it is necessary to prepare a corresponding initial quantum state, often requiring the generation of complex quantum states such as entangled states. Known methods for this include generating Dicke states [7]. In this work, we focus on this aspect and propose a more efficient and flexible method for generating initial quantum states.
In this study, we propose an approach called "Two-Step QAOA." This method follows the conventional design approach of formulating constraints and cost functions but separates these terms into two distinct quantum circuit stages rather than computing them simultaneously within the same equation. By doing so, it transforms soft constraints into hard constraints, simplifies the creation of initial conditions, and enables more efficient optimization.
In the first stage, the structure of the QUBO formulation is analyzed to distinguish one-hot constraints from other terms. The one-hot constraints are then processed separately and converted into hard constraints and their corresponding quantum states. In the second stage, the hard constraints derived from the one-hot constraints are used to create a new QUBO formulation, which is then optimized more effectively using QAOA.
@@
The objective of this study is to utilize the Two-Step QAOA approach to reduce the complexity of setting initial conditions and improve the overall performance of QAOA. This method is expected to pave the way for solving complex societal problems such as logistics, resource allocation, and scheduling. Furthermore, as the quantum state generation process is heuristic, various improvements, including approximations, can be anticipated.
@@ 3.2 Initialization
The QAOA protocol begins with the preparation of an initial state, denoted as $|\psi_{0}\rangle$. This initial state isn’t chosen arbitrarily - its selection is carefully determined according to the specific Hamiltonian being used in the algorithm. Specifically, it is prepared as an eigenstate of the mixer Hamiltonian. We’ll delve into the mathematical details and representation of this state in the following chapter.
@@ 3.3 Application of Cost and Mixer Hamiltonians
3.4 Time Evolution
The overall time evolution process of QAOA is mathematically represented as follows. Starting from the initial state, the cost and mixer Hamiltonians are applied $p$ times:
@@ 4.1 Mixer Hamiltonian
In standard QAOA, the mixer Hamiltonian $H_{M}$ is typically constructed using the Pauli X operator, expressed as:
$H_{M}=\sum_{j=1}^{n}X_{j}$ (6)
where $X_{j}$ represents the Pauli X operator acting on qubit j.
However, in the Quantum Alternating Operator Ansatz, more general mixer Hamiltonians can be constructed using different combinations of Pauli operators. For example, the XY mixer Hamiltonian can be written as:
$H_{M}=\frac{1}{2}\sum_{i,j}(X_{i}X_{j}+Y_{i}Y_{j})$ (7)
This flexibility in choosing mixer Hamiltonians $H_{M}$ allows for exploration of different mixing strategies that may be more suitable for specific optimization problems and can potentially lead to better performance compared to using just the standard $X$ Hamiltonian.
@@ 4.2 Initialization
When using standard Pauli X mixer Hamiltonians, we prepare simple superposition states as initial quantum states. The equation is as follows:
$|\psi_{0}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{z\in\{0,1\}^{n}}|z\rangle$ (8)
For more sophisticated mixer Hamiltonians constructed from combinations of Pauli operators, we need to prepare more complex quantum states that include not only superposition but also quantum entanglement as eigenstates. Here is an example of an initial quantum state corresponding to the XY mixer: Dicke states, which can be prepared as an entangled state where k qubits out of n qubits are in the excited state (|1⟩).
$|\psi_{0}\rangle=|D(n,k)\rangle=\frac{1}{\sqrt{C(n,k)}}\sum_{P}P(|1\rangle^{\otimes k}|0\rangle^{\otimes(n-k)})$ (9)
where:
  • $C(n,k)$ represents the binomial coefficient
  • $P$ denotes the sum over all possible permutations
  • $n$ is the total number of qubits
  • $k$ is the number of excited states
Preparing Dicke states exactly in quantum circuits can often lead to circuit expansion and become a challenging problem.
5 Cost and Constraint Terms in QUBO Formulation
@@
In QAOA, constructing the cost Hamiltonian involves creating terms that minimize the objective function and satisfy necessary constraints. These terms are expressed in the form of a Quadratic Unconstrained Binary Optimization (QUBO) problem, which is widely used in optimization tasks.
The QUBO formulation is defined as:
$\min_{\mathbf{x}}\;\mathbf{x}^{T}Q\mathbf{x},$ (10)
where $\mathbf{x}$ is a vector of binary variables ($x_{i}\in\{0,1\}$), and $Q$ is a symmetric matrix that encodes the problem. Each element $Q_{ij}$ represents the weight of the quadratic term involving $x_{i}$ and $x_{j}$, while diagonal elements $Q_{ii}$ represent the linear coefficients of $x_{i}$.
5.1 Cost Terms
The cost terms in the QUBO formulation represent the objective function that needs to be minimized. For example, in a graph-based optimization problem such as Maximum Cut (MaxCut), the cost Hamiltonian can be written as:
$H_{\text{cost}}=\sum_{(i,j)\in E}w_{ij}x_{i}(1-x_{j})$ (11)
where:
  • $E$ is the set of edges in the graph,
  • $w_{ij}$ is the weight of the edge between vertices $i$ and $j$,
  • $x_{i}$ and $x_{j}$ are binary variables indicating the assignment of vertices to two subsets.
In matrix form, the QUBO representation of this cost term is:
$H_{\text{cost}}=\mathbf{x}^{T}Q_{\text{cost}}\mathbf{x}$ (12)
where the matrix $Q_{\text{cost}}$ encodes the edge weights and their contributions to the objective function.
5.2 Constraint Terms
In many problems, constraints are essential to ensure that solutions are valid. These constraints are incorporated into the QUBO formulation by introducing penalty terms. A common approach is to add quadratic penalties that enforce the constraints.
For example, consider a constraint that the sum of binary variables must equal a specific value $k$:
$\sum_{i}x_{i}=k$ (13)
This constraint can be enforced in the QUBO formulation by adding a penalty term:
$H_{\text{constraint}}=\lambda\left(\sum_{i}x_{i}-k\right)^{2}$ (14)
where $\lambda$ is a tuning parameter that controls the strength of the penalty.
In matrix form, this can also be expressed as:
$H_{\text{constraint}}=\mathbf{x}^{T}Q_{\text{constraint}}\mathbf{x}$ (15)
where $Q_{\text{constraint}}$ contains the coefficients derived from the penalty function.
5.3 Balancing Cost and Constraint Terms
The final QUBO Hamiltonian combines the cost and constraint terms:
$H_{\text{QUBO}}=H_{\text{cost}}+H_{\text{constraint}}$ (16)
In practice, tuning parameters like $\lambda$ are critical for balancing the trade-off between minimizing the objective function and satisfying the constraints. Improper tuning can lead to solutions that either violate constraints or fail to optimize the objective function effectively.
@@ 5.3 Balancing Cost and Constraint Terms
Determining the appropriate value for $\lambda$ often requires iterative optimization. This process involves adjusting $\lambda$ iteratively until a satisfactory balance is achieved, making sure that the final solution both minimizes the cost and satisfies the constraints. The search space expands with the addition of constraint terms, making it harder to find the global minimum.
6 Method: A Two-Step Initialization Approach
For certain constraint conditions, there is a way to avoid calculating the balance between such troublesome cost functions and constraints. For k-hot constraints where k qubits among n qubits are in state |1> while the rest are in |0>, by preparing appropriate initial quantum states and computing through XY mixer exploration, the constraints can be excluded from the cost Hamiltonian as hard constraints and instead incorporated into the mixer Hamiltonian.
However, preparing Dicke states, which correspond to XY mixer initial quantum states, typically involves difficulties. While there are exact rules for preparing quantum states, there are concerns about excessive quantum circuit overhead when preparing these states on early quantum computers. In this proposal, to address these issues, a two-stage approach is used where QAOA is heuristically utilized to easily generate initial quantum states corresponding to the XY mixer.
@@ 6.2 Construct Mixer Hamiltonian
In this method, among the computational steps that typically consist of initial quantum state preparation and QAOA step execution, we also utilize QAOA steps to create the initial quantum state. Therefore, we use different Hamiltonians in two separate QAOA executions.
@@
$H_{M_{2}}=H_{XY}=\sum_{i,j}(X_{i}X_{j}+Y_{i}Y_{j})+\ldots$ (19)
@@ 6.3 Modify Cost Hamiltonian
Next, we will adjust the cost function. Since we will run QAOA twice, we need to prepare two cost Hamiltonians. In the proposed method, preparing the cost Hamiltonians is straightforward.
For the first QAOA, the cost Hamiltonian $H_{C_{1}}$ is prepared by removing the adjustment variables from the constraint conditions. Specifically, it becomes:
$H_{C_{1}}=H_{constraint^{\prime}}=\left(\sum_{i}x_{i}-k\right)^{2}$ (20)
For the second QAOA, we set up a cost function that removes constraint conditions from the cost Hamiltonian.
$H_{C_{2}}=H_{cost}$ (21)
We have prepared two combinations of cost Hamiltonians and mixer Hamiltonians for executing QAOA twice. Finally, we will verify the calculation steps.
6.4 Procedure Overview of a two-step approach
First, we start from a general superposition state and execute the first QAOA to obtain an intermediate quantum state.
@@ 6.4 Procedure Overview of a two-step approach
The cost Hamiltonian used in this first QAOA is the constraint condition without adjustment variables. When this QAOA is executed, it generates a quantum state that satisfies the constraint conditions. This is expected to be close to the initial quantum state of the XY mixer.
Next, we use the intermediate quantum state created in the first QAOA as the initial state for the second QAOA. Since this quantum state is expected to be close to the initial quantum state corresponding to the XY mixer, when the final calculation is performed, we obtain results that satisfy the k-hot constraint while calculating the cost Hamiltonian.
In this step, since adjustment variables do not appear in the final formulation, there is no need to balance the cost function and constraint conditions. Optimize the parameters $\gamma_{1},\gamma_{2},\beta_{1},\beta_{2}$ through classical optimization techniques to find the optimal solution.
@@ 7 Discussion
Even without detailed knowledge of complex quantum circuit construction methods for precise Dicke state generation, this approach offers a significant advantage: as long as soft constraints can be expressed in equation form, they can be utilized to generate states approximating Dicke states. This accessibility makes the method particularly valuable for practical applications.